# Graphon Quantum Filtering systems

### v1

@@ 2 Preliminaries
In this section we recall main notions from quantum filtering theory, graphon theory, and the Fubini extension.
@@ 2.1 Quantum Filtering
As stated in Section 1, the framework for continuous observation of a quantum system is provided by quantum filtering theory and its fundamental principle of non-demolition measurement, pioneered by Belavkin in [10, 11, 12]. This leads to a diffusive stochastic evolution of quantum state in the space of density operators
$\mathrm{d}\rho_{t}$ $=-\mathrm{i}[\tilde{H},\rho_{t}]\mathrm{d}t+\left(L\rho_{t}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\rho_{t}\}\right)\mathrm{d}t+\sqrt{\eta}\left(L\rho_{t}+\rho_{t}L^{\dagger}-\mathrm{t}\mathrm{r}\big{(}(L+L^{\dagger})\rho_{t}\big{)}\rho_{t}\right)\mathrm{d}W_{t},$
$\rho_{0}$ $\in\mathcal{S}(\mathbb{H}),$ (2.1)
where $(W_{t})_{t\geq 0}$ is a standard Wiener process, $\tilde{H}=\tilde{H}^{\dagger}\in\mathcal{B}(\mathbb{H})$ is the Hamiltonian of the system, $L\in\mathcal{B}(\mathbb{H})$ is the measurement operator. The scalar $\eta\in(0,1]$ denotes the measurement efficiency. The corresponding measurement records associated with the operator $L$ is given by
@@
In the filtering approach the process $(\rho_{t})_{t\geq 0}$ is our best estimation of the system’s quantum state given the measurements records $(Y_{s})_{0\leq s\leq t}$.
The following proposition deal with important remark of wave-function representation of the state dynamics when purity is conserved.
Proposition 1.
If the measurement efficiency is perfect i.e $\eta=1$, and the initial state is pure $(\rho_{0}=\ket{\psi_{0}}\bra{\psi_{0}})$, then the purity is preserved and the dynamic (2.2) has pure state representation as follow,
$\mathrm{d}\psi_{t}$ $=-\Big{(}\mathrm{i}\Big{(}\tilde{H}-\langle\frac{L+L^{\dagger}}{2}\rangle_{\psi_{t}}\frac{L-L^{\dagger}}{2\mathrm{i}}\Big{)}+\frac{1}{2}\Big{(}(L-\langle\frac{L+L^{\dagger}}{2}\rangle_{\psi_{t}})^{\dagger}(L-\langle\frac{L+L^{\dagger}}{2}\rangle_{\psi_{t}})\Big{)}\Big{)}\psi_{t}\mathrm{d}t$
$\qquad\quad+\Big{(}L-\langle\frac{L+L^{\dagger}}{2}\rangle_{\psi_{t}}\Big{)}\psi_{t}\mathrm{d}W_{t},$
or in equivalent way,
$\mathrm{d}\psi_{t}$ $=-\Big{(}\mathrm{i}\tilde{H}+\frac{1}{2}L^{\dagger}L-\frac{\langle L+L^{\dagger}\rangle_{\psi_{t}}}{2}L+\frac{\langle L+L^{\dagger}\rangle_{\psi_{t}}^{2}}{8}\Big{)}\psi_{t}\mathrm{d}t+\Big{(}L-\langle\frac{L+L^{\dagger}}{2}\rangle_{\psi_{t}}\Big{)}\psi_{t}\mathrm{d}W_{t}.$
The equation written in pure state takes usually the name of stochastic Schrodinger equation.
Remark 1.
@@
The Belavkin equation may also describe a more general setting regarding simultaneous continuous observation of several physical observables corresponding to a family of measurement operators $(L_{1},\dots,L_{N})$ and $N$-dimensional Wiener process $\mathbf{W}=(W^{1},\dots,W^{N})$. The equation takes the following form
$\mathrm{d}\rho_{t}=$ $-\mathrm{i}[\tilde{H},\rho_{t}]\mathrm{d}t+\sum_{l=1}^{N}\left(L_{l}\rho_{t}L_{l}^{\dagger}-\frac{1}{2}\{L^{\dagger}_{l}L_{l},\rho_{t}\}\right)\mathrm{d}t$
$+\sqrt{\eta}\sum_{l=1}^{N}\left(L_{l}\rho_{t}+\rho_{t}L_{l}^{\dagger}-\mathrm{t}\mathrm{r}\big{(}(L_{l}+L^{\dagger}_{l})\rho_{t}\big{)}\rho_{t}\right)\mathrm{d}W^{l}_{t}.$
@@
The main interest for these models is that they allow for the consideration of quantum feedback control through feedback (Markovian feedback in the terminology of Mean-field games). In fact, as in stochastic control with partial information, the separation theorem [17] allows us to treat the problem of estimating the state from the optimal control problem. The theory of quantum filtering then reduces the analysis to the controlled version of (2.2), where the Hamiltonian part is modified by allowing control parameter $\alpha\in\mathcal{U}$ such that the equation becomes (see e.g., [70, 32])
$\mathrm{d}\rho_{t}$ $=-\mathrm{i}[\tilde{H}+\alpha_{t}\hat{H},\rho_{t}]\mathrm{d}t+\left(L\rho_{t}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\rho_{t}\}\right)\mathrm{d}t+\sqrt{\eta}\left(L\rho_{t}+\rho_{t}L^{\dagger}-\mathrm{t}\mathrm{r}\big{(}(L+L^{\dagger})\rho_{t}\big{)}\rho_{t}\right)\mathrm{d}W_{t}.$ (2.2)
@@
We can now state the problem of optimal control given $\alpha$, by fixing an objective function
$\mathcal{J}((\alpha_{s})_{0\leq s\leq T},\rho_{0}):=\mathbb{E}\Big{[}\int_{0}^{T}C(\alpha_{s},\rho_{s})\mathrm{d}s+F(\rho_{T})\Big{]},$
@@
where $C$ and $F$, representing respectively current and terminal costs, are assumed to be bounded and differentiable. The objective is to find $\alpha^{\star}$ that minimizes $\mathcal{J}$. The natural question of the Hamilton-Jacobi-Bellman equation is treated in [34, 15, 14].
@@ 2.2 Graphon theory
We give a brief summary of the notion of graphon theory relevant to our work, see [50] for more details.
Let $I:=[0,1]$. The set $I$ is an index set labeling a continuum of particles. The Lebesgue space over $I$ is denoted by $(I,\mathcal{B}_{I},\lambda_{I})$. Denote by $\mathcal{W}$ the space of all bounded symmetric measurable functions $w:I^{2}\rightarrow\mathbb{R}$. The elements of $\mathcal{W}$ will be called kernels. Let $\mathcal{W}_{0}$ be the subset of $\mathcal{W}$ such that $0\leq w\leq 1$, where elements of $\mathcal{W}_{0}$ will be called graphons.
An important notion in graphon theory is that of the operator associated to any kernel. For a graphon $w$ we can define the operator $\mathcal{T}_{w}:L^{\infty}(I)\to L^{1}(I)$ as follows: for any $\phi\in L^{\infty}(I)$,
$\mathcal{T}_{w}(\phi)(u):=\int_{0}^{1}w(u,v)\phi(v)\mathrm{d}v.$
@@
$\|w\|_{\square}:=\sup_{\mathrm{B}_{1},\mathrm{B}_{2}\in\mathcal{B}_{I}}\Big{|}\int_{\mathrm{B}_{1}\times\mathrm{B}_{2}}w(u,v)\mathrm{d}u\mathrm{d}v\Big{|},$
@@
$\|\mathcal{T}_{w}\|_{\text{op}}:=\sup_{\|\phi\|_{\infty}\leq 1}\int_{I}\Big{|}\int_{I}w(u,v)\phi(v)\mathrm{d}v\Big{|}\mathrm{d}u.$
@@
$\|w\|_{L^{1}}$ $:={\int_{0}^{1}\int_{0}^{1}|w(u,v)|\mathrm{d}v\mathrm{d}u}.$
According to [50, lemma 8.11], for every kernel $w\in\mathcal{W}$, we have equivalence between operator norm and cut norm,
$\|w\|_{\square}\leq\|\mathcal{T}_{w}\|_{\text{op}}\leq 4\|w\|_{\square}.$ (2.3)
@@
Definition 1 (Step Kernel).
Given a graph $G_{N}=(V_{N},E_{N})$, one can define a step function $w^{G_{N}}:[0,1]^{2}\rightarrow\mathbb{R}$ that associates with the adjacency matrix $((\xi_{pq}^{N})_{p,q\in V_{N}}$. This association is constructed by partitioning the interval $[0,1]$ into subintervals $I_{p}=\left(p-\frac{1}{N},p\right]$ for $p=\frac{1}{N},\frac{2}{N},\dots,\frac{N}{N}$, such that
@@
This definition of step kernel enables us to consider the limiting behavior of a sequence of graphs using the cut metric. A sequence of dense graphs $(G_{N})_{N\geq 0}$ is said to converge to a graphon if their corresponding step kernels converge (in cut norm) to a graphon function $w$.
@@ 2.3 Fubini extension theory
In order to overcome a serious measurability issue that arises when constructing a continuum family of pairwise independents stochastic processes, the notion of Fubini extension was introduced in [66], indeed one may wonder if pairwise independents family $(X^{u})_{u\in I}$ can be constructed on probability spaces $(\Omega,\mathcal{F},\mathbb{Q})$ so that the process $X\colon I\times\Omega\ni(u,\omega)\mapsto X^{u}(\omega)$ satisfies joint measurability property. To do that we shall construct such processes on the extensions of the product space $(I\times\Omega,\mathcal{B}_{\mathcal{I}}\otimes\mathcal{F},\lambda_{\mathcal{I}}\otimes\mathbb{Q})$, which are called Fubini’s extensions.
We follow here the same approach as in [3, 28, 29] to define graphon SDE in the appropriate Fubini extension space and recall main ingredients for stochastic processes on Fubini extended space, see e.g., [23, Section 3.7] for more details.
Definition 2 (Pairwise independence).
@@
Let $(I,\mathcal{B}_{\mathcal{I}},\lambda_{\mathcal{I}})$ and $(\Omega,\mathcal{F},\mathbb{Q})$ be two probability spaces. A process $X:I\times\Omega\to S$, where $S$ is a Polish space, is said to be essentially pairwise independent (e.p.i.) if for $\lambda$-almost every $u\in I$, the random variables $X^{u}$ and $X^{v}$ are independent for $\lambda$-almost every $v\in I$.
Definition 3 (Fubini extension [29]).
@@
Let $(I,\mathcal{B}_{\mathcal{I}},\lambda_{I})$ be a finite measure space and $(\Omega,\mathcal{F},\mathbb{Q})$ a probability space. A triple $(I\times\Omega,\mathcal{V},\mathcal{Q})$ is called a Fubini extension of the product space $(I\times\Omega,\mathcal{B}_{\mathcal{I}}\otimes\mathcal{F},\lambda_{I}\otimes\mathbb{Q})$ if:
@@
  • the functions $g_{u}:\omega\to g(u,\omega)$ and $g_{\omega}:u\to g(u,\omega)$ are integrable on $(\Omega,\mathcal{F},\mathbb{Q})$ for $\lambda-$a.e. $u\in I$, and on $(I,\mathcal{I},\lambda)$ for $\mathbb{Q}$-a.e. $\omega\in\Omega$, respectively.
@@
  • the integrals $\int_{\Omega}g_{u}(\omega)\mathbb{Q}(\mathrm{d}\omega)$ and $\int_{I}g_{\omega}(u)\lambda(\mathrm{d}u)$ are integrable, respectively, on $(I,\mathcal{I},\lambda)$ and $(\Omega,\mathcal{F},\mathbb{Q})$, with
@@
    $\int_{I\times\Omega}g(\omega,u)\mathcal{Q}(\mathrm{d}u,\mathrm{d}\omega)$ $=\int_{I}\Bigl{(}\int_{\Omega}g_{u}(\omega)\mathbb{Q}(\mathrm{d}\omega)\Bigr{)}\lambda(\mathrm{d}u)=\int_{\Omega}\Bigl{(}\int_{I}g_{\omega}(u)\lambda(\mathrm{d}u)\Bigr{)}\mathbb{Q}(\mathrm{d}\omega).$
We denote $(I\times\Omega,\mathcal{I}\boxtimes\mathcal{F},\lambda\boxtimes\mathbb{Q}):=(I\times\Omega,\mathcal{V},\mathcal{Q}).$
@@
The following theorem from [67] guarantees the existence of Fubini extension space carrying a collection of e.p.i. jointly measurable random variables.
Theorem 1.
@@
There exists a probability space $(I,\mathcal{I},\lambda)$ extending $(I,\mathcal{B}_{I},\lambda_{I})$, a probability space $(\Omega,\mathcal{F},\mathbb{Q})$, and a Fubini extension $(I\times\Omega,\mathcal{I}\boxtimes\mathcal{F},\lambda\boxtimes\mathbb{Q})$ of $(I\times\Omega,\mathcal{I}\otimes\mathcal{F},\lambda\otimes\mathbb{Q})$ such that for any measurable mapping
$\varphi\colon(I,\mathcal{I},\lambda)\to\mathcal{P}(S),$
@@
the set of Borel probability measures on $S$, there exists an $\mathcal{F}\,\boxtimes\,\mathcal{I}$-measurable process
$X\colon I\times\Omega\to S$
such that the random variables $X^{u}(\bullet)$ are e.p.i. and
$\mathbb{Q}\circ(X^{u})^{-1}=\varphi(u)\quad\text{for all }u\in I.$
Given a Polish space $S$, let $\mathcal{C}:=\mathcal{C}([0,T],S)$ space of continuous from $[0,T]$ to $S$, we denote $L^{2}_{\mathcal{C}}(I\times\Omega,\mathcal{Q})$ the space of square integrable processes with respect to the box product defined above. That is,
$L^{2}_{\mathcal{C}}(I\times\Omega,\mathcal{Q})=\Bigl{\{}X:I\times\Omega\to C,\quad\mathcal{V}-\text{measurable}\quad\text{and}\quad\int_{I\times\Omega}\|X^{u}(\omega)\|_{\mathcal{C}}^{2}\mathcal{Q}(\mathrm{d}u,\mathrm{d}\omega)<\infty\Bigr{\}}.$

### v3

@@ 2 Preliminaries
In this section we recall main notions and results from quantum filtering theory, graphon theory, and the Fubini extension.
@@ 2.1 Quantum filtering
The theory of quantum filtering, with its fundamental principle of non-demolition measurement, was pioneered by Belavkin [12, 13, 14]. It provides a rigorous framework for the real-time observation and control of quantum dynamical systems without destroying their state. For a comprehensive introduction to the theory, we refer the reader to [21, 22, 36].
In the case of homodyne measurement, the evolution of the state is described by a diffusive stochastic differential equation, known as the Belavkin equation:
$\mathrm{d}\rho_{t}$ $=-\mathrm{i}[\tilde{H},\rho_{t}]\mathrm{d}t+\left(L\rho_{t}L^{\dagger}-\tfrac{1}{2}\{L^{\dagger}L,\rho_{t}\}\right)\mathrm{d}t+\sqrt{\eta}\left(L\rho_{t}+\rho_{t}L^{\dagger}-\mathrm{t}\mathrm{r}\big((L+L^{\dagger})\rho_{t}\big)\rho_{t}\right)\mathrm{d}W_{t},$
$\rho_{0}$ $\in\mathcal{S}(\mathbb{H}),$ (1)
[…]
$Y_{t}$ $=W_{t}+\sqrt{\eta}\int_{0}^{t}\mathrm{tr}\big((L+L^{\dagger})\rho_{s}\big)\mathrm{d}s.$
@@
In the filtering approach, the process $(\rho_{t})_{t\geq 0}$ represents the best estimate of the system’s quantum state given the measurement record $(Y_{s})_{0\leq s\leq t}$.
Remark 2.1.
@@
The Belavkin equation can also be formulated in a more general setting involving the simultaneous continuous observation of several physical observables, corresponding to a family of measurement operators $(L_{1},\dots,L_{M})$ and an $M$-dimensional Wiener process ${W}=(W^{1},\dots,W^{M})$. In this case, the equation takes the form
$\mathrm{d}\rho_{t}=$ $-\mathrm{i}[\tilde{H},\rho_{t}]\mathrm{d}t+\sum_{l=1}^{M}\left(L_{l}\rho_{t}L_{l}^{\dagger}-\frac{1}{2}\{L^{\dagger}_{l}L_{l},\rho_{t}\}\right)\mathrm{d}t$
$+\sqrt{\eta}\sum_{l=1}^{M}\left(L_{l}\rho_{t}+\rho_{t}L_{l}^{\dagger}-\mathrm{t}\mathrm{r}\big((L_{l}+L^{\dagger}_{l})\rho_{t}\big)\rho_{t}\right)\mathrm{d}W^{l}_{t}.$
@@
The main interest of these models lies in the fact that they allow for the incorporation of quantum feedback control (Markovian feedback in the terminology of mean-field games). Indeed, as in stochastic control with partial information, the separation theorem [19] makes it possible to decouple the state estimation problem from the optimal control problem. The theory of quantum filtering then reduces the analysis to the controlled version of (2), where the Hamiltonian part is modified by introducing a real control parameter $\alpha$ in the set of admissible controls $\mathcal{U}$; so that the equation becomes (see, e.g., [79, 33]):
$\mathrm{d}\rho_{t}$ $=-\mathrm{i}[\tilde{H}+\alpha_{t}\hat{H},\rho_{t}]\mathrm{d}t+(L\rho_{t}L^{\dagger}-\tfrac{1}{2}\{L^{\dagger}L,\rho_{t}\})\mathrm{d}t+\sqrt{\eta}(L\rho_{t}+\rho_{t}L^{\dagger}-\mathrm{tr}\!\big((L+L^{\dagger})\rho_{t}\big)\rho_{t})\mathrm{d}W_{t},$ (2)
where $\hat{H}=\hat{H}^{\dagger}\in\mathcal{B}(\mathbb{H})$ is the control Hamiltonian.
@@
We can now formulate the optimal control problem given $\alpha$ by fixing an objective function
$\mathcal{J}\Big((\alpha_{s})_{0\leq s\leq T},\rho_{0}\Big):=\mathbb{E}_{\mathbb{P}}\Big[\int_{0}^{T}C(\alpha_{s},\rho_{s})\mathrm{d}s+F(\rho_{T})\Big],$
@@
where $C$ and $F$, representing respectively the running and terminal costs, are assumed to be bounded and differentiable. The objective is to find $\alpha^{\star}$ that minimizes $\mathcal{J}$. The associated Hamilton–Jacobi–Bellman equation has been studied in [37, 17, 16].
The following proposition addresses an important point concerning the wave-function representation of the state dynamics.
Proposition 2.2.
[…]
$\mathrm{d}\psi_{t}$ $=-\Big(\mathrm{i}\Big(\tilde{H}-\langle\tfrac{L+L^{\dagger}}{2}\rangle_{\psi_{t}}\tfrac{L-L^{\dagger}}{2\mathrm{i}}\Big)+\tfrac{1}{2}\Big((L-\langle\tfrac{L+L^{\dagger}}{2}\rangle_{\psi_{t}})^{\dagger}(L-\langle\tfrac{L+L^{\dagger}}{2}\rangle_{\psi_{t}})\Big)\Big)\psi_{t}\mathrm{d}t$
$\qquad\quad+\Big(L-\langle\tfrac{L+L^{\dagger}}{2}\rangle_{\psi_{t}}\Big)\psi_{t}\mathrm{d}W_{t},$ (3)
or, equivalently,
$\mathrm{d}\psi_{t}$ $=-\Big(\mathrm{i}\tilde{H}+\tfrac{1}{2}L^{\dagger}L-\tfrac{\langle L+L^{\dagger}\rangle_{\psi_{t}}}{2}L+\tfrac{\langle L+L^{\dagger}\rangle_{\psi_{t}}^{2}}{8}\Big)\psi_{t}\mathrm{d}t+\Big(L-\langle\tfrac{L+L^{\dagger}}{2}\rangle_{\psi_{t}}\Big)\psi_{t}\mathrm{d}W_{t}.$
This pure-state formulation is commonly referred to as the stochastic Schrödinger equation.
Proof.
Let’s denote the drift (resp. diffusive) part of the equation (2.2) by $\mathrm{F}_{t}$ (resp. $\mathrm{G}_{t}$). Then by Itô’s rule,
[…]
and,
[…]
where the commutator with $\tilde{H}$ comes from the Hamiltonian part of $\mathrm{F}_{t}$. Since $\tilde{H}$ is self-adjoint, we have
[…]
and,
[…]
$|.\displaystyle\quad+\tfrac{1}{4}\langle L+L^{\dagger}\rangle_{\psi_{t}}^{2}\ket{\psi_{t}}\bra{\psi_{t}}$.
Adding the two contributions gives
[…]
and for the diffusion term we have
[…]
Therefore, denoting $|\rho_{t}=\ket{\psi_{t}}\bra{\psi_{t}}$, the process satisfies
$\mathrm{d}\rho_{t}$ $=-\mathrm{i}[\tilde{H},\rho_{t}]\,\mathrm{d}t+\Big(L\rho_{t}L^{\dagger}-\tfrac{1}{2}\{L^{\dagger}L,\rho_{t}\}\Big)\mathrm{d}t+\Big(L\rho_{t}+\rho_{t}L^{\dagger}-\langle L+L^{\dagger}\rangle_{\psi_{t}}\rho_{t}\Big)\mathrm{d}W_{t},$
where
$\langle L+L^{\dagger}\rangle_{\psi_{t}}=\mathrm{t}\mathrm{r}\!\left((L+L^{\dagger})\rho_{t}\right).$
This completes the proof. ∎
@@ 2.2 Graphon theory
We give a brief summary of the notion of graphon theory relevant to our work, see [55] for further details.
Let $I:=[0,1]$. The set $I$ is an index set labeling a continuum of particles. The Lebesgue space over $I$ is denoted by $(I,\mathfrak{B}_{I},\lambda_{I})$ where $\lambda_{I}(\mathrm{d}u):=\mathrm{d}u$ is the Lebesgue measure. Denote by $\mathcal{W}$ the space of all bounded symmetric measurable functions $w:I^{2}\rightarrow\mathbb{R}$. The elements of $\mathcal{W}$ will be called kernels.
Let $\mathcal{W}_{0}$ be the subset of $\mathcal{W}$ such that $0\leq w\leq 1$, where elements of $\mathcal{W}_{0}$ will be called graphons. Intuitively, graphons generalize graphs to a continuum of vertices: $w(u,v)$ represents the weight of the edge between $u$ and $v$. These objects are ideal for studying convergent sequences of dense graphs.
For a graphon $w$ we can define the operator $\mathcal{T}_{w}:L^{\infty}(I)\to L^{1}(I)$ as follows: for any $\phi\in L^{\infty}(I)$,
$\mathcal{T}_{w}(\phi)(u):=\int_{I}w(u,v)\phi(v)\mathrm{d}v.$
@@
$\|w\|_{\square}:=\sup_{\mathrm{B}_{1},\mathrm{B}_{2}\in\mathfrak{B}_{I}}\Big|\int_{\mathrm{B}_{1}\times\mathrm{B}_{2}}w(u,v)\mathrm{d}u\mathrm{d}v\Big|,$
@@
$\|\mathcal{T}_{w}\|_{\text{op}}:=\sup_{\|\phi\|\leq 1}\int_{I}\Big|\int_{I}w(u,v)\phi(v)\mathrm{d}v\Big|\mathrm{d}u.$
@@
$\|w\|_{L^{1}}$ $:={\int_{I}\int_{I}|w(u,v)|\mathrm{d}v\mathrm{d}u}.$
According to [55, lemma 8.11], for every kernel $w\in\mathcal{W}$, the operator norm and the cut norm are equivalent,
$\|w\|_{\square}\leq\|\mathcal{T}_{w}\|_{\text{op}}\leq 4\|w\|_{\square}.$ (4)
@@
Definition 2.3 (Step Kernel).
Given a finite graph $G_{N}=(V_{N},E_{N})$, where $V_{N}=\{\tfrac{1}{N},\dots,\tfrac{N}{N}\}$ with adjacency matrix $(\xi_{pq}^{N})_{p,q\in V_{N}}$, the associated step kernel $w^{G_{N}}:[0,1]^{2}\rightarrow\mathbb{R}$ is defined by partitioning the unit interval $[0,1]$ into subintervals $I_{p}=\left(p-\frac{1}{N},p\right]$ for $p\in V_{N}$, such that
@@
This definition of a step kernel allows us to study the limiting behavior of a sequence of graphs using the cut metric. A sequence of dense graphs $(G_{N})_{N\geq 0}$ is said to converge to a graphon $w$ if their associated step kernels converge to $w$ in the cut norm.
@@ 2.3 Fubini extension theory
To resolve the measurability difficulties that arises when constructing a continuum family of pairwise independents stochastic processes, the notion of Fubini extension was introduced in [73]. Indeed, one may ask whether a pairwise independent family $\Theta:=\Big\{\theta_{u}:u\in I\Big\}$ can be constructed on a probability space $(\Omega,\mathcal{F},\mathbb{Q})$ such that the process $\Theta\colon I\times\Omega\ni(u,\omega)\mapsto\Theta(u,\omega)=\theta_{u}(\omega)$ satisfies joint measurability property. To achieve this, such processes must be constructed on extensions of the standard product space $(I\times\Omega,\mathfrak{B}_{I}\otimes\mathcal{F},\lambda_{\mathcal{I}}\otimes\mathbb{Q})$.
We follow a similar approach to [5, 30, 31, 2] to define stochastic processes on the appropriate Fubini extension space, see e.g., [25, Section 3.7] for an introduction to the formalism of Fubini extensions. We begin by recalling an important notion due to Sun [72] which shows that all the notions of independence are, in fact, almost identical to their pairwise counterparts in an ideal setting of a continuum of random variables.
Definition 2.4 (Essential pairwise independence).
@@
Let $(I^{\prime},\mathcal{I}^{\prime},\lambda^{\prime})$ and $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{Q}^{\prime})$ be two probability spaces, and let $E$ be a Banach space . A random variable $\Theta:I^{\prime}\times\Omega^{\prime}\to E$, is said to be essentially pairwise independent (e.p.i.) if for $\lambda^{\prime}$-almost every $u\in I^{\prime}$, the random variables $\Theta(u,.)$ and $\Theta(v,.)$ are independent for $\lambda^{\prime}$-almost every $v\in I^{\prime}$.
We now turn to the notion of a Fubini extension, which provides the appropriate probability space for hosting a jointly measurable e.p.i. family. This extension enlarges the classical product space in such a way that both independence and Fubini’s theorem are preserved, and it serves as the foundational framework for all the results developed in this paper.
Definition 2.5 (Fubini extension).
@@
Let $(I^{\prime},\mathcal{I}^{\prime},\lambda^{\prime})$ and $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{Q}^{\prime})$ be two probability spaces. A triple $(I^{\prime}\times\Omega^{\prime},\mathcal{V}^{\prime},\mathcal{Q}^{\prime})$ is called a Fubini extension of the product space $(I^{\prime}\times\Omega^{\prime},\mathcal{I}^{\prime}\otimes\mathcal{F}^{\prime},\lambda^{\prime}\otimes\mathbb{Q}^{\prime})$ if for any Banach space $E$-valued $\mathcal{Q}^{\prime}$-integrable function $g$ on $(I^{\prime}\times\Omega^{\prime},\mathcal{V}^{\prime})$ :
@@
  • the functions $g_{u}:\omega\mapsto g(u,\omega)$ and $g_{\omega}:u\mapsto g(u,\omega)$ are integrable on $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{Q}^{\prime})$ for $\lambda^{\prime}$-a.e. $u\in I^{\prime}$, and on $(I^{\prime},\mathcal{I}^{\prime},\lambda^{\prime})$ for $\mathbb{Q}^{\prime}$-a.e. $\omega\in\Omega^{\prime}$, respectively.
@@
  • the functions $u\mapsto\int_{\Omega^{\prime}}g_{u}(\omega)\mathbb{Q}^{\prime}(\mathrm{d}\omega)$ and $\omega\mapsto\int_{I^{\prime}}g_{\omega}(u)\lambda^{\prime}(\mathrm{d}u)$ are integrable, respectively, on $(I^{\prime},\mathcal{I}^{\prime},\lambda^{\prime})$ and $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{Q}^{\prime})$, with
@@
    $\int_{I^{\prime}\times\Omega^{\prime}}g(u,\omega)\mathcal{Q}^{\prime}(\mathrm{d}u,\mathrm{d}\omega)$ $=\int_{I^{\prime}}\Bigl(\int_{\Omega^{\prime}}g_{u}(\omega)\mathbb{Q}^{\prime}(\mathrm{d}\omega)\Bigr)\lambda^{\prime}(\mathrm{d}u)=\int_{\Omega^{\prime}}\Bigl(\int_{I^{\prime}}g_{\omega}(u)\lambda^{\prime}(\mathrm{d}u)\Bigr)\mathbb{Q}^{\prime}(\mathrm{d}\omega).$
Although it is common in the literature to denote the Fubini extension $(I^{\prime}\times\Omega^{\prime},\mathcal{V}^{\prime},\mathcal{Q}^{\prime})$ by $(I^{\prime}\times\Omega^{\prime},\mathcal{I}^{\prime}\boxtimes\mathcal{F}^{\prime},\lambda^{\prime}\boxtimes\mathbb{Q}^{\prime})$, we avoid this notation here for clarity.
@@
The following result from [74] guarantees the existence of Fubini extension space carrying a collection of e.p.i. jointly measurable random variables.
Theorem 2.6.
@@
[74, Theorem $1$] There exists a probability space $(I,\mathcal{I},\lambda)$ extending $(I,\mathfrak{B}_{I},\lambda_{I})$, a probability space $(\Omega,\mathcal{F},\mathbb{Q})$, and a Fubini extension $(I\times\Omega,\mathcal{V},\mathcal{Q})$ of the usual product space $(I\times\Omega,\mathcal{I}\otimes\mathcal{F},\lambda\otimes\mathbb{Q})$ such that for any measurable mapping
$\varphi:I\to\mathrm{P}(S),$
@@
where $S$ is a Polish space and $\mathrm{P}(S)$ denotes the set of Borel probability measures on $S$, there exists a $\mathcal{V}$-measurable process,
$\Theta:I\times\Omega\to S,$
such that the random variables $\Theta(u,.)$ are e.p.i. and
$\mathbb{Q}\circ(\Theta({u},.)^{-1})=\varphi(u),\quad\text{for all }u\in I.$
Throughout this paper, we work on the space $(I,\mathcal{I},\lambda)$ for the index space.
Now we state an important result on the Fubini extension, the Exact Law of Large Numbers (ELLN), which allows us to replace the stochastic mean by a deterministic mean.
Proposition 2.7.
[73, Corollary 2.10] Let $\Theta$ be an integrable random variable on $(I\times\Omega,\mathcal{V},\mathcal{Q})$. If $\Theta$ is essentially pairwise independent, then the sample mean is the same as the mean of $\Theta$,
$\int_{I}\Theta(u,\tilde{\omega})\lambda(\mathrm{d}u)=\int_{I\times\Omega}\Theta(u,\omega)\mathcal{Q}(\mathrm{d}u,\mathrm{d}\omega),\;\;\forall\tilde{\omega}\in\Omega,\;\mathbb{Q}\text{-a.s}$
We next state a version of Girsanov’s theorem, adapted to a family of e.p.i. Wiener processes in the Fubini extension setting, as detailed in our companion paper [2].
Proposition 2.8.
[2, Theorem 4] Let $(I\times\Omega,\mathcal{V},\mathcal{Q})$ be a Fubini extension probability space. Consider a collection of e.p.i. Wiener processes $\mathbf{B}=\big\{B^{u}:u\in I\big\}$, and a real valued process $(\Theta_{t})_{t\geq 0}=\Big\{(\theta_{t,u})_{t\geq 0}:u\in I\Big\}$, such that for all $(t,u)\in[0,T]\times I$, $\theta_{t,u}$ is $\sigma(B_{s}^{u};s\leq t)$-measurable.
Define the process $(\mathcal{E}_{t})_{t\geq 0}=\Big\{(\mathcal{E}_{t,u})_{t\geq 0}:u\in I\Big\}$ by
$\mathcal{E}_{t}:=\exp\Big\{\int_{0}^{t}\Theta_{s}\mathrm{d}\mathbf{B}_{s}-\frac{1}{2}\int_{0}^{t}\Theta_{s}^{2}\mathrm{d}s\Big\}.$
[…]
$\tilde{\mathbf{B}}_{t}$ $:=\mathbf{B}_{t}-\int_{0}^{t}\Theta_{s}\mathrm{d}s.$